I sent this problem in but no one helped me so here it is again:
You order items from a menu that total $7.85. Your bill comes to $8.30, including the tax. What is the percent of the tax? Round to the nearest tenth. After you do the problem estimate a 15% tip for this order before the tax.

Thanks whoever does this a lot!!!!!!

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

8.30-7.85=$.45 amount of tax
.45/7.85=.057 or .06 or 6% for the tax amount.
7.85*.15=$1.18 for the tip on the base price.

Answer by edjones(8007)

let x=1+decimal of the tax.
7.85x=8.3
7.85x/7.85=8.3/7.85
x= 1.057
tax=(1.057-1)*100
=.057*100
5.7%
.
To calculate a 15% tip on 8.30:
move the decimal point one to the left (.83) and add 1/2 the result to the result. .83+.41=1.24 est.
Be a big spender and give $1.25
Ed